A) Find the LCD, covert each rational expression into an equivalent expression with the LCD as the denominator.
4/x-4 , 5/6-a


B)If the given expressions were used as denominators of rational expressions, then what would be the LCD for each group of denominators? See example 3
3/4b^2-0 2b/2b+3 -5/2b^2-3b


Any help with the above two would greatly appreciated. I am trying to better understand how to do these and my text book is of no help.

Answer by stanbon(48568)

A) Find the LCD, covert each rational expression into an equivalent expression with the LCD as the denominator.
4/x-4 , 5/6-a
lcd = (x-4)(6-a) because those denominators go not share a factor.
Converting you should divide the given denominator into the lcd
and multiply the numerator by the quotient:
For example: lcd/(x-4) = (6-a)
1st fraction becomes [4(6-a)/lcd]
-------------------------------
Then, lcd/(6-a)= (x-4)
2nd fraction becomes 5(x-4)/lcd
=======================================


B)If the given expressions were used as denominators of rational expressions, then what would be the LCD for each group of denominators? See example 3
3/4b^2-0 2b/2b+3 -5/2b^2-3b
-------
Comment I think you have a typo in the denominator of the 1st fraction.
------------
Factor each expression as follows:
3/4b^2 , 2b/(2b+3), -5/[b(2b-3)]
-----------
The lcd must contain each factor in its highest power:
lcd = 4b^2(2b+3)(2b-3)
===============================
Cheers,
Stan H.